# Words with factor complexity $2n+1$ and minimal critical exponent

### v2

@@ 7 A sharpened structure theorem and open problems
$(\gamma^{R})^{\infty}(2)=20102102010\cdots.$
Then the factors of $\mathbf{G}$$^{\prime}$ are precisely the reverses of the factors of $\mathbf{G}$}. One can easily adjust the proof of the Morphism Lemma to prove
@@
Suppose that ${\mathbf{z}}\in\{0,1,2\}^{\omega}$ is low. Suppose that a final segment of $\mathbf{z}$} has the form $h({\mathbf{u}})$, some ${\mathbf{u}}\in\{0,1,2\}^{\omega}$ where $h=[a,b,c]$, some $a,b,c\in\Sigma_{3}^{+}$ such that:
@@
    Word $b$ is a suffix of $a$, which is a suffix of $c$.
@@
    Words $b$ and $c$ have a common prefix $p$ such that $|p|\geq|b|/2$.
@@
Let ${\mathbf{w}}\in\Sigma_{3}^{\omega}$ be low, and let $\mathbf{w}$} have factor complexity $2n+1$. One of the following holds:
  • For some non-erasing morphism $h$ in Figure 1, $h(g)$ is a factor of $\mathbf{w}$} whenever $g$ is a factor of $\mathbf{G}$};
  • For some non-erasing morphism $h$ in Figure 2, $h(g)$ is a factor of $\mathbf{w}$} whenever $g$ is a factor of $\mathbf{G}$$^{\prime}$.
@@
We note in passing that the third morphism in Figure 1, namely
$h=[021201212,0212012,021201212012],$
discovered by the computer implementation of the proof, is a conjugate of
@@
Morphism $h$ could thus be replaced in the figure/proof by the identity morphism. However, we did not complicate our proof implementation by searching for simplifications and/or duplications of morphisms.
The sharpened theorem says that the factor set of $\mathbf{w}$} is the same as the factor set of $h({\mathbf{G}})$ for a morphism in Figure 1, or $h({\mathbf{G}}^{\prime})$ for a morphism in Figure 2. It is not known for which, if any, of the $h$ the words $h({\mathbf{G}})$ and $h({\mathbf{G}}^{\prime})$ have factor complexity $2n+1$, and what the critical exponents of these words are. We make the following conjecture:
@@
Appendix 1: ‘Human readable’ proof of the resolution of the case $S=\{01,12,20,00\}$ of the Structure Theorem
Our Python implementation outputs the following proof:
We assume $F$ includes
$\{01,12,20,00\}$
@@
Word w must omit a factor from
$\{0210,0211,2210,2211\}$
@@
$\{02110,02210,02211,10211,10221,11022,21021,21102,22102,22110\}$
Word w must omit a factor from
$\{11021,21022\}$
@@
  Case 1.1: $F$ includes
$\{01,12,20,00,0210,11021\}$
@@
$\{02110221022,02110221102,02210211022,02210221102,02211022102,$
10211022102,10211022110,10221021102,10221022110,10221102210,10211022102,10211022110,10221021102,10221022110,10221102210,
11022102110,11022102211,11022110221,21021102210,21021102211,11022102110,11022102211,11022110221,21021102210,21021102211,
21022110221,21102210211,21102210221,21102211022,21022110221,21102210211,21102210221,21102211022,
22102110221,22102211022,22110221021,22110221022}22102110221,22102211022,22110221021,22110221022\}
In this case w must omit factor 02110221021
@@
Author’s comment: It turns out that when we find a length-250 low word $w$ omitting the current list of missing factors, the complexity is $2n+1$ for $n=1,2,\ldots,20$. We therefore (in step C.a) search for a morphism, considering factors $b$ of $w$ of length up to 3, and finding the return words of $b$.
The return words of 210 are among
$\{2102,21021102,210221102,2102110221102\}$
Word 2102 cannot occur twice in a concatenation of these
Author’s comment: If 2102 occurs twice, it occurs in the context $s2102t$, where $s,t$ are two of the return words. However, one checks that each such word $s2102t$ contains an $r$ power, $r\geq 5/2$.
Thus a final segment of w is concatenated from
$\{21021102,210221102,2102110221102\}$
Taking conjugates, a final segment is the image under morphism
@@
which satisfies the Dual Morphism Lemma.
Author’s comment: Since $\mathbf{w}$} has a final segment of the form $h({\mathbf{u}})$, the Dual Morphism Lemma shows that $\mathbf{w}$} has critical exponent at least $\mu$, and the current subcase is resolved.
Case 1.2: $F$ includes
$\{01,12,20,00,0210,21022\}$
The return words of 210 are:
$\{21021102,2102110221102,21021102211021102,2102110221102110221102\}$
Word 21021102 cannot occur twice in a concatenation of these.
    Thus a final segment of w is concatenated from:
$\{2102110221102,21021102211021102,2102110221102110221102\}$
Taking conjugates, a final segment is the image under morphism
[21021102211021102,2102110221102,2102110221102110221102][21021102211021102,2102110221102,2102110221102110221102]
@@
    Case 2: $F$ includes
$\{01,12,20,00,0211\}$
@@
$\{02102,02210,02211,10210,10221,11021,11022,21021,21022,21102,22102,22110\}$
Word w has $12>2\times 5+1$ factors of length 5 . This is impossible.
@@
$\{02102,02110,02211,10210,10211,10221,11021,11022,21021,21022,21102,22110\}$
Word w has $12>2\times 5+1$ factors of length 5 . This is impossible.
  Case 4: $F$ includes
$\{01,12,20,00,2211\}$
@@
$\{02102,02110,02210,10210,10211,10221,11021,11022,21021,21022,21102,22102\}$
Word w has $12>2\times 5+1$ factors of length 5 . This is impossible.
@@
Author’s comment: The proof in the present case is much shorter than for the resolution of $S=\{11,22\}$. In that longer resolution the computer frequently invokes step A of the algorithm (not used in the present case) avoid repetition of case analysis.
Appendix 2: Python implementation of the case resolution software
The following Python worksheet performs all the calculations mentioned in the paper:
import math
import time
def complexityBreak(w):
# Let the set of factors of w of length n be c[n]. If for some
# 0<i<depth+1 we have |c[i]|>2i+1, return c[i] for the least such
# i. Otherwise return 0.
    depth=20
    for i in range(1,depth+1):
        c=[]
        for j in range(len(w)-i+1):
            if not(w[j:j+i]in c):
                c.append(w[j:j+i])
        if (len(c)>2*i+1):
            c.sort()
            return(c) #
    return(0)
def goodMorphism(L):
# Does L=[a,b,c] satisfy the conditions of the Morphism Lemma?
    def prefix(a,b):
    # Does word b have prefix a?
        if (len(b)<len(a)):
            return(False)
        else:
            return(b[:len(a)]==a)
    def goodsuff(a,c):
    # Do c and a have a common suffix of length |a|/2?
        for i in range(math.ceil(len(a)/2),len(a)+1):
            if (a[-i:]==c[-i:]):
                return(True)
        return(False)
    a=L[0]
    b=L[1]
    c=L[2]
    return(prefix(a,b)and prefix(b,c)and goodsuff(a,c)
           and (2*len(a)>=len(c)))
def dualMorphism(L):
# Does L=[a,b,c] satisfy the conditions of the Dual Morphism Lemma?
    A=L[0][::-1]
    B=L[1][::-1]
    C=L[2][::-1]
    L=[A,B,C]
    return(goodMorphism(L))
def morph(L):
# Returns a conjugate of morphism L which satisfies the Morphism Lemma
# or the Dual Morphism Lemma. If no such conjugate exists, returns 0.
    def morphism_conjugates(L):
    # Given a morphism L (a list of length 3) such that L[0]<=L[1]<=L[2],
    # returns a list of the conjugates of L.
        conjugates=[]
        conjugates.append(L)
        M=L
    # To begin, M is a copy of L. We iteratively conjugate by (prefix)
    # letters
        i=0
        while(i<len(L[0])):
            if ((M[0][0]==M[1][0])and(M[1][0]==M[2][0])):
                # M = [xa,xb,xc], so B=[ax,bx,cx] is a conjugate.
                B=[]
                for j in range(3):
                    B.append(M[j][1:]+M[j][0])
                M=B
                conjugates.append(B)
            i=i+1
    # Again, M is a copy of L. We iteratively conjugate by (suffix)
    # letters
        M=L
        i=1
        while(i<len(L[0])):
            if ((M[0][-1]==M[1][-1])and(M[1][-1]==M[2][-1])):
                # M = [ax,bx,cx], so B=[xa,xb,xc] is a conjugate.
                B=[]
                for j in range(3):
                    B.append(M[j][-1]+M[j][:-1])
                M=B
                conjugates.append(B)
            i=i+1
        return(conjugates)
    C=morphism_conjugates(L)
    for M in C:
        if goodMorphism(M):
            N=[M[1],M[0],M[2]]
            return([1,N])
        if dualMorphism(M):
            N=[M[1],M[0],M[2]]
            return([2,N])
    return([0,[]])
def fhpf(w): # Word w has no suffix of exponent 5/2 or greater
    p=1 #potential period of suffix of exponent 5/2 or greater
    while (5*p<=2*len(w)):
        if (w[-3*p//2:]==w[-5*p//2:-p]):
            return(False)
        p=p+1
    return(True)
def test(w,forbidden_factors):
# This returns true if no suffix of w is in forbidden_factors or has
# exponent at least 5/2
    for f in forbidden_factors:
        if (len(w)>=len(f)):
            if (w[-len(f):]==f):
                return(False)
    return(fhpf(w))
def backtrack(target,forbidden_factors):
# Call a ternary word _low if it contains no factor having exponent
# 5/2 or greater. This routine returns a low word w starting
# with 0 of length target such that w contains no factor in F. If no
# such word exists, it returns the length of the longest low word
# starting with 0 with no factor in forbidden_factors.
    largest_letter=’2’
    w=’0’
    mx=1 # Greatest length attained so far
    while(1==1):
        success=test(w,forbidden_factors)
        if (success):
            if (len(w)>mx):
                mx=len(w)
            if (len(w)==target):
                return(w)
            else:
                w=w+’0’
        else:
            while((len(w)>1) and w[-1]==largest_letter):
                w=w[:-1]
            if (len(w)==1):
#               Search fails. Longest word has length mx.
                return(mx)
            else:
                w=w[:-1]+chr(ord(w[-1])+1)
def good(w,forbidden_factors):
# Call a word _good if it is low and has no factor in
# forbidden_factors.
    for i in range(1,len(w)+1):
        if not(test((w[:i]),forbidden_factors)):
            return(False)
    return(True)
def doubler(n,forbidden_factors):
# This returns the set of all good words of length n. It produces
# the words recursively, testing concatenations of good words of
# lengths ceil(n/2) and floor(n/2).
    if (n==0):
        return([’’])
    if (n==1):
        return([’0’,’1’,’2’])
    G=[]
    for s in doubler(math.ceil(n/2),forbidden_factors):
        for t in doubler(n//2,forbidden_factors):
            u=s+t
            if good(s+t,forbidden_factors):
                G.append(s+t)
    return(G)
def returns(b,depth,forbidden_factors):
# Finds certain return words of b: words bu such that bub is
# good, |u|<=depth, and there are exactly two instances of b in bub,
# namely as prefix and suffix.
    myReturns=[]
    for i in range(depth+1):
        forbidden_factors.append(b) # We seek good words u not
                                    # containing b
        U=doubler(i,forbidden_factors)
        forbidden_factors.pop()
        for u in U:
            if good(b+u+b,forbidden_factors):
                if ((b+u+b)[1:].find(b)==(len(b)+len(u)-1)):
                    myReturns.append(b+u)
    return(myReturns)
def factors(w,n):
# Returns the set of factors of w of length n
    f=[]
    for j in range(len(w)-n+1):
        if not(w[j:j+n] in f):
            f.append(w[j:j+n])
    f.sort()
    return(f)
Morphisms=[]
Dual_Morphisms=[]
longest_b=1
def findMorph(w,k,forbidden_factors):
# Given word w and set of excluded words forbidden_factors, we
# consider factors b # of w of length up to k. Suppose that every
# good word of length v+1 contains b. We find the return words bu.
# If there are exactly three return words (or four return words
# of which the first can only occur once) then any good word has
# a final segment concatenated from the three return words. We
# form a morphism from these return words. We then test whether
# one of its conjugates satisfies the Morphism Lemma or the Dual
# Morphism Lemma,
    global Morphisms
    global Dual_Morphisms
    global longest_b
    n=1
    while(n<=k):
        Fact=factors(w,n)
        for b in Fact:
            forbidden_factors.append(b)
            depth=backtrack(200,forbidden_factors)
            forbidden_factors.pop()
            if(isinstance(depth,int)):
                # Longest good word not containing b has length
                # depth
                Returns=returns(b,depth,forbidden_factors)
                if(len(Returns)>=3):
                    if(len(Returns)<=4):
                            myMorph=[Returns[-3],Returns[-2],Returns[-1]]
                            N=morph(myMorph)
                            if(N[0]==1):
                            # This conjugate of myMorph satisfies
                            # the Morphism Lemma. We check that
                            # Returns[0] cannot appear twice
                            # in a good concatenation of the return
                            # words
                                check=True
                                for p in Returns:
                                    for s in Returns:
                                        if (good(p+Returns[0]+s,
                                                 forbidden_factors)):
                                            check=False
                                if (check):
                                    print(indent+’The return words ’+
                                          ’of ’,b,’ are:’+’\\’+’\\’)
                                    print(Returns,’\\’+’\\’)
                                    print(indent+’Word ’,Returns[0],
                                    ’cannot occur twice in a ’+
                                    ’concatenation of these.\\’+’\\’)
                                    print(indent+’Thus a final ’+
                                    ’segment of w is concatenated ’+
                                    ’from:’,’\\’+’\\’)
                                    print(myMorph,’\\’+’\\’)
                                    print(indent+’Taking conjugates,’+
                                    ’ a final segment is the image ’+
                                    ’under morphism’,’\\’+’\\’)
                                    print(N[1],’\\’+’\\’)
                                    if (not(N[1]in Morphisms)):
                                        Morphisms.append(N[1])
                                    print(indent+’which satisfies’+
                                    ’ the Morphism Lemma.\\’+’\\’)
                                    return()
                            if(N[0]==2): # This conjugate of myMorph
                                         # satisfies the Dual Morphism Lemma
                                check=True
                                for p in Returns:
                                    for s in Returns:
                                        if (good(p+Returns[0]+s,
                                                 forbidden_factors)):
                                            check=False
                                if (check):
                                    print(indent+’The return words’+
                                    ’ of ’,b,’ are among ’,
                                    Returns,’\\’+’\\’)
                                    print(indent+’Word ’,Returns[0],
                                    ’cannot occur twice in a ’+
                                    ’concatenation of these’,’\\’+
                                    print(indent+’Thus a final ’
                                    +’segment of w is ’+
                                    ’ concatenated from ’,myMorph,
                                    print(indent+’Taking conjugates,’+
                                    ’ a final segment is the image ’+
                                    ’under morphism’,’\\’+’\\’)
                                    print(N[1],’\\’+’\\’)
                                    if (not(N[1]in Dual_Morphisms)):
                                        Dual_Morphisms.append(N[1])
                                    print(indent+’which satisfies’+
                                    ’ the Dual Morphism ’+
                                    ’Lemma.’,’\\’+’\\’)
                                    return()
        n=n+1
        longest_b=max(longest_b,n)
    print(indent+’No morphism found up to length ’, k,’\\’+’\\’)
    return()
indent = ’’
def resolveCase(forbidden_factors,caseString):
    global indent
    global resolved_cases
    global resolved_case_labels
    indent=’~’
    for i in range(len(caseString)):
        indent=indent+’~’
    k=3 # Maximum length of b considered in return words bu
    if (caseString==’’):
        print(indent+’We assume F includes ’,forbidden_factors,
    w=backtrack(250,forbidden_factors)
    if (isinstance(w,int)):
        print(’No good word longer than ’,w,’ avoids these factors.’
        F=[]
        for f in forbidden_factors:
            F.append(f)
        resolved_cases.append(F)
        resolved_case_labels.append(caseString)
        print(resolved_cases)
        print(resolved_case_labels)
        return()
#   The case being resolved is labelled by forbidden_factors.
#   However, we may find that additional factors are also necessarily
#   avoided in this case. We will sharpen our further backtracking
# searches within this case by avoiding these additional factors.
#   To put it another way, when there is exactly one unneeded factor,
#   we would get one new subcase. Since there is no branching, it is
#   more intuitive to simply say the subcase is the "same" case, but
#   we specify additional avoided factors in factorsToAvoid. The "base
#   case" is still labelled by the set forbidden_factors
    factorsToAvoid=[]
    for f in forbidden_factors:
        factorsToAvoid.append(f)
    unneeded_factors=[]
    while(len(unneeded_factors)<2):
        w=backtrack(250,factorsToAvoid)
        N=complexityBreak(w)
        if (N==0): #complexity is good
            M=findMorph(w,k,factorsToAvoid)
            F=[]
            for f in forbidden_factors:
                F.append(f)
            resolved_cases.append(F)
            resolved_case_labels.append(caseString)
            return()
        else:
            Mx=0
            needed_factors=[]
            unneeded_factors=[]
            for s in N:
                factorsToAvoid.append(s)
                w=backtrack(250,factorsToAvoid)
                factorsToAvoid.pop()
                if (isinstance(w,int)):
                    Mx=max(Mx,w)
                    needed_factors.append(s)
                else:
                    unneeded_factors.append(s)
            if(len(needed_factors)>0):
                print(indent+’Every good word longer than ’,Mx,
                ’must include factor(s)’,’\\’+’\\’)
                print(needed_factors,’\\’+’\\’)
                if(len(needed_factors)>2*len(N[0])+1):
                    print(indent+’Word w has ’,len(needed_factors),
                    ’> 2 x’,len(N[0]),’+ 1 factors of length ’,
                    len(N[0]) ,’. This is impossible.’,’\\’+’\\’)
                    F=[]
                    for f in forbidden_factors:
                        F.append(f)
                    resolved_cases.append(F)
                    resolved_case_labels.append(caseString)
                    return()
            if (len(unneeded_factors)==1):
                print(indent+’In this case w must omit factor ’,
                unneeded_factors[0],
                s=unneeded_factors[0]
                factorsToAvoid.append(s)
    unneeded_factors.sort()
    j=0
    print(indent+’Word w must omit a factor from’,’\\’+’\\’)
    print(unneeded_factors,’\\’+’\\’)
    print(indent+’This gives rise to ’,len(unneeded_factors),
    ’ cases.’,’\\’+’\\’)
    for s in unneeded_factors:
        if (caseString==’’):
            subcase=chr(ord(’1’)+j)
        else:
            subcase=caseString+’.’+chr(j+49)
        j=j+1
        factorsToAvoid.append(s)
        print(indent+’Case ’+subcase+’: F includes ’,factorsToAvoid,
        # Test whether case was resolved previously
        resolved=False
        for c in resolved_cases:
            resolved=True
            for d in c:
                if not(d in factorsToAvoid):
                    resolved=False
                    break
            if (resolved):
                i=resolved_cases.index(c)
                print(’This was previously resolved in Case ’+
                resolved_case_labels[i]+’.’,’\\’+’\\’)
                break
        if(not(resolved)):
            resolveCase(factorsToAvoid,subcase)
            factorsToAvoid.pop()
    F=[]
    for f in forbidden_factors:
        F.append(f)
    resolved_cases.append(F)
    resolved_case_labels.append(caseString)
    return()
print(’Here is the resolution of S={22,11}:’)
print(’ ’)
resolved_cases=[]
resolved_case_labels=[]
startTime=time.time()
resolveCase([’22’,’11’],’’)
endTime=time.time()
print(’Computation took ’,(endTime-startTime),’ seconds.’)
print(’ ’)
print(’ ’)
print(’ ’)
print(’Here is the resolution of S={01,12,20,00}:’)
print(’ ’)
resolved_cases=[]
resolved_case_labels=[]
startTime=time.time()
resolveCase([’01’,’12’,’20’,’$0$’],’’)
endTime=time.time()
print(’Computation took ’,(endTime-startTime),’ seconds.’)
print(’ ’)
print(’ ’)
print(’ ’)
print(’Here are the morphisms satisfying the Morphism Lemma used ’+
’in the case resolutions:’)
print(’ ’)
for m in Morphisms:
    print(m)
print(’ ’)
print(’ ’)
print(’ ’)
print(’Here are the morphisms satisfying the Dual Morphism Lemma ’+
’used in the case resolutions:’)
print(’ ’)
for m in Dual_Morphisms:
    print(m)
print(’ ’)
print(’ ’)
print(’ ’)
print(’Here is the length of the longest word in the backtracks of’+
’ Lemma 2:’)
S=[’1121’, ’1122’, ’1211’, ’1212’, ’1221’, ’2112’,
’2121’, ’2122’, ’2211’, ’2212’]
F=[’10’,’20’,’$0$’]
longest=0
for s in S:
        for t in S:
            if (s!=t):
                F.append(s)
                F.append(t)
                w=backtrack(500,F)
                longest=max(w,longest)
                F.pop()
                F.pop()
print(’ ’)
print(longest)

### v3

@@ 7 A sharpened structure theorem and open problems
${\mathbf{G}}^{\prime}=(\gamma^{R})^{\infty}(2)=20102102010\cdots.$
(An anonymous referee points out that ${\mathbf{G}}^{\prime}=\pi({\mathbf{p}})$ where $\mathbf{p}$} is the sequence studied in [11] and $\pi$ is the permutation of $[2,0,1]$ of $\Sigma_{3}$. See the entry for sequence A287072 in the On-Line Encyclopedia of Integer Sequences.) Then the factors of $\mathbf{G}$$^{\prime}$ are precisely the reverses of the factors of $\mathbf{G}$}. One can easily adjust the proof of the morphism lemma to prove
@@
Suppose that ${\mathbf{z}}\in\Sigma_{3}^{\omega}$ is low. Suppose that a suffix of $\mathbf{z}$} has the form $h({\mathbf{u}})$, some ${\mathbf{u}}\in\Sigma_{3}^{\omega}$ where $h=[a,b,c]$, some $a,b,c\in\Sigma_{3}^{+}$ such that
@@
    The word $b$ is a suffix of $a$, which is a suffix of $c$.
@@
    The words $b$ and $c$ have a common prefix $p$ such that $|p|\geq|b|/2$.
@@
Let ${\mathbf{w}}\in\Sigma_{3}^{\omega}$ be eligible. One of the following holds.
  • For some non-erasing morphism $h$ in Figure 1 and a permutation $\pi$ of $\Sigma_{3}$, the word $\pi(h(g))$ is a factor of $\mathbf{w}$} whenever $g$ is a factor of $\mathbf{G}$};
  • For some non-erasing morphism $h$ in Figure 2 and some permutation $\pi$ of $\Sigma_{3}$, the word $\pi(h(g))$ is a factor of $\mathbf{w}$} whenever $g$ is a factor of $\mathbf{G}$$^{\prime}$.
@@
The morphisms in Figures 1 and 2 were output by the computer implementation of our proof. In fact, the implementation gave 26 different morphisms, each produced by applying a permutation to one of the listed morphisms. Some further simplification is possible. For example, consider the third morphism in Figure 1, namely
$h=[021201212,0212012,021201212012].$
This is a conjugate of
@@
Morphism $h$ could thus be replaced in Figure 1 by the identity morphism. We did not complicate our proof implementation by searching for simplifications/duplications of morphisms.
Open Problem 1.
Give a version of the sharpened structure theorem with an ‘aesthetically optimal’ list of morphisms.
The sharpened theorem says that the set of factors of $\mathbf{w}$} is the same as the set of factors of $h({\mathbf{G}})$ for a morphism in Figure 1, or $h({\mathbf{G}}^{\prime})$ for a morphism in Figure 2. It is not known for which, if any, of the $h$ the words $h({\mathbf{G}})$ and $h({\mathbf{G}}^{\prime})$ have factor complexity $2n+1$, and what the critical exponents of these words are. We make the following conjecture:
@@
Taken together with Carpi and De Luca’s result [5] for Sturmian words, we have the beginnings of a Dejean’s Theorem for Arnoux-Rauzy words; we know the minimal critical exponent in the cases $k=2$ and $k=3$.
Open Problem 2.
For each integer $k\geq 4$, give an Arnoux-Rauzy word over $\Sigma_{k}$ with the least critical exponent.
Appendix: ‘Human readable’ proof of the resolution of the case $S=\{01,12,20,00\}$ of the structure theorem
The Python implementation outputs the following proof:
We assume $F$ includes [01, 12, 20, 00].
@@
The word w must omit a factor from
[0210, 0211, 2210, 2211].
@@
Case 1: $F$ includes [01, 12, 20, 00, 0210]
Case 2: $F$ includes [01, 12, 20, 00, 0211]
Case 3: $F$ includes [01, 12, 20, 00, 2210]
Case 4: $F$ includes [01, 12, 20, 00, 2211]
@@
[02110, 02210, 02211, 10211, 10221, 11022, 21021, 21102, 22102, 22110].
The word w must omit a factor from
[11021, 21022].
@@
Case 1.1: $F$ includes [01, 12, 20, 00, 0210, 11021]
Case 1.2: $F$ includes [01, 12, 20, 00, 0210, 21022]
Case 1.1: $F$ includes [01, 12, 20, 00, 0210, 11021].
@@
[02110221022, 02110221102, 02210211022, 02210221102, 02211022102, 10211022102, 10211022110, 10221021102, 10221022110, 10221102210, 11022102110, 11022102211, 11022110221, 21021102210, 21021102211, 21022110221, 21102210211, 21102210221, 21102211022, 22102110221, 22102211022, 22110221021, 22110221022].
In this case w must omit the factor 02110221021.
@@
The $210$-blocks are among [2102, 21021102, 210221102, 2102110221102].
Author’s comment: It turns out that when we find a length-250 low word $w$ omitting the current list of missing factors, the complexity is $2n+1$ for $n=1,2,\ldots,20$. We therefore, in step C.a, search for a morphism, considering factors $b$ of $w$ with $|b|\leq 3$ and finding the $b$-blocks of $w$.
The word 2102 cannot occur twice in a concatenation of these
Author’s comment: If 2102 occurs twice, it occurs in the context $s2102t$, where $s,t$ are two of the $210$-blocks. However, one checks that each such word $s2102t$ contains an $r$ power with $r\geq 5/2$.
Thus a suffix of w is concatenated from [21021102, 210221102, 2102110221102].
Taking conjugates, a suffix is the image under morphism
[221102210, 21102210, 2110221102210].
which satisfies the dual morphism lemma.
(This concludes Case 1.1).
Author’s comment: Since $\mathbf{w}$} has a suffix of the form $h({\mathbf{u}})$, the dual morphism lemma shows that $\mathbf{w}$} has critical exponent at least $\mu$, and the current subcase is resolved.
Case 1.2: $F$ includes [01, 12, 20, 00, 0210, 21022].
The 210-blocks are
[21021102, 2102110221102, 21021102211021102, 2102110221102110221102].
The word 21021102 cannot occur twice in a concatenation of these.
Thus a suffix of w is concatenated from
@@
Taking conjugates, a suffix is the image under morphism
[21021102211021102, 2102110221102, 2102110221102110221102].
@@
(This concludes Case 1.2).
(This concludes Case 1).
Case 2: $F$ includes [01, 12, 20, 00, 0211].
@@
[02102, 02210, 02211, 10210, 10221, 11021, 11022, 21021, 21022, 21102, 22102, 22110].
The word w has $12>2\times 5+1$ factors of length 5. This is impossible.
(This concludes Case 2).
@@
[02102, 02110, 02211, 10210, 10211, 10221, 11021, 11022, 21021, 21022, 21102, 22110].
The word w has $12>2\times 5+1$ factors of length 5. This is impossible.
(This concludes Case 3).
Case 4: $F$ includes [01, 12, 20, 00, 2211].
@@
[02102, 02110, 02210, 10210, 10211, 10221, 11021, 11022, 21021, 21022, 21102, 22102].
The word w has $12>2\times 5+1$ factors of length 5. This is impossible.
(This concludes Case 4).
@@
Author’s comment: The proof in the present case is much shorter than for the resolution of $S=\{11,22\}$. In that longer resolution the computer frequently invokes step A of the algorithm (not used in the present case) to avoid repeating case analysis which was already carried out.